### v1

@@ Abstract
Unsigned distance functions (UDFs) have been a vital representation for open surfaces. With different differentiable renderers, current methods are able to train neural networks to infer a UDF by minimizing the rendering errors on the UDF to the multi-view ground truth. However, these differentiable renderers are mainly handcrafted, which makes them either biased on ray-surface intersections, or sensitive to unsigned distance outliers, or not scalable to large scale scenes. To resolve these issues, we present a novel differentiable renderer to infer UDFs more accurately. Instead of using handcrafted equations, our differentiable renderer is a neural network which is pre-trained in a data-driven manner. It learns how to render unsigned distances into depth images, leading to a prior knowledge, dubbed volume rendering priors. To infer a UDF for an unseen scene from multiple RGB images, we generalize the learned volume rendering priors to map inferred unsigned distances in alpha blending for RGB image rendering. Our results show that the learned volume rendering priors are unbiased, robust, scalable, 3D aware, and more importantly, easy to learn. We evaluate our method on both widely used benchmarks and real scenes, and report superior performance over the state-of-the-art methods. Project page: https://wen-yuan-zhang.github.io/VolumeRenderingPriors/.
Keywords:
Unsigned distance function Volume rendering Implicit reconstruction
Figure 1: We highlight our multi-view reconstruction results from UDFs learned on real-captured open surface scenes and indoor scenes. The two sides of a surface are colored in white and beige, respectively. Comparing with NeuS [47] and the state-of-the-art UDF reconstruction method NeUDF [29], our method does not produce artifacts and recovers more accurate and smooth geometries on both open and closed surfaces.
1 Introduction
Neural implicit representations have become a dominated representation in 3D computer vision. Using coordinate based deep neural networks, a mapping from locations to attributes at these locations like geometry [43, 39], color [10, 38], and motion [15] can be learned as an implicit representation. Signed distance function (SDF) [39] and unsigned distance function (UDF) [9] are widely used implicit representations to represent either closed surfaces [27, 22, 8] or open surfaces [58, 17]. We can learn SDFs or UDFs from supervisions like ground truth signed or unsigned distances [4, 39], 3D point clouds [32, 46, 7, 60, 34, 33, 59] or multi-view images [52, 30, 20, 56, 62]. Compared to SDFs, it is a more challenging task to estimate a UDF due to the sign ambiguity and the boundary effect, especially under a multi-view setting.
@@ 1 Introduction
Recent methods [30, 29, 13, 35] mainly infer UDFs from multi-view images through volume rendering. Using different differentiable renderers, they can render a UDF into RGB or depth images which can be directly supervised by the ground truth images. These differentiable renderers are mainly handcrafted equations which are either biased on ray-surface intersections, or sensitive to unsigned distance outliers, or not scalable to large scale scenes. These issues make them struggle with recovering accurate geometry. Fig. 23 details these issues by comparing error maps on depth images rendered by different differentiable renderers. We render the ground truth UDF into depth images using different renderers from 100 different view angles, and report the average rendering error on each one of 55 shapes that are randomly sampled from each one of the 55 categories in ShapeNet [5] in Fig. 2. Using the latest differentiable renderers from NeuS-UDF [47] (using UDF as input to NeuS), NeUDF [29] and NeuralUDF [30], the rendered depth images and their error maps in Fig. 3 (a) to (c) show that these issues cause large errors even using the ground truth UDF as inputs. Therefore, how to design better differentiable renderers for UDF inference from multi-view images is still a challenge.
@@
To resolve these issues, we introduce a novel differentiable renderer for UDF inference from multi-view images through volume rendering. Instead of handcrafted equations used by the latest methods [30, 29, 13, 35], we employ a neural network to learn to become a differentiable renderer in a data-driven manner. Using UDFs and depth images obtained from meshes as ground truth, we train the neural network to map a set of unsigned distances at consecutive locations along a ray into weights for alpha blending, so that we can render depth images, and produce the rendering errors to the ground truth as a loss. We make the neural network observe different variations of unsigned distance fields during training, and learn the knowledge of volume rendering with unsigned distances by minimizing the rendering loss. The knowledge we call volume rendering prior is highly generalizable to infer UDFs from multi-view RGB images in unobserved scenes. During testing, we use the pre-trained network as a differentiable renderer for alpha blending. It renders unsigned distances inferred by a UDF network into RGB images which can be supervised by the observed RGB images. Our results in Fig. 2 and Fig. 3 (e) show that we produce the smallest rendering errors among all differentiable renderers for UDFs, which is even more accurate than NeuS-SDF [47] (rendering with ground truth SDF) in Fig. 3 (d). Extensive experiments in our evaluations show that the learned volume rendering priors are unbiased, robust, scalable, 3D aware, and more importantly, easy to learn. We conduct evaluations in both widely used benchmarks and real scenes, and report superior performance over the state-of-the-art methods. Our contributions are listed below,
@@
2 Related Work
@@
Multi-view 3D reconstruction. Multi-view 3D reconstruction aims to reconstruct 3D shapes from calibrated images captured from overlapping viewpoints. The key idea is to leverage the consistency of features across different views to infer the geometry. MVSNet [51] is the first to introduce the learning-based idea into traditional MVS methods. Following studies explore the potential of MVSNet in different aspects, such as training speed [54, 49], memory consumption [50, 16], network structure [14] and generalization [57]. These techniques produce depth maps or 3D point clouds. To obtain meshes as final 3D representations, additional procedures such as TSDF-fusion [11] or classic surface reconstruction [23] methods are used, which is complex and not intuitive.
@@
Learning SDFs from Multi-view Images. Instead of 3D point clouds estimated by MVS methods, recent methods [47, 52, 48] directly estimate SDFs through volume rendering from multi-view images for continuous surface representations. The widely used strategy is to render the estimated SDF into RGB images [12, 40, 28] or depth images [55, 45, 3] which can be supervised by the ground truth images. The key to make the whole procedure differentiable is various differentiable renderers [47, 3, 37] which transform signed distances into weights for alpha blending during rendering. Some methods modify the rendering equations to use more 2D supervisions like normal maps [44], detected planes [18], and segmentation maps [25] to pursue higher reconstruction efficiency. However, the SDFs that these methods aim to learn are only for closed surfaces, which is limited to represent open surfaces.
@@
Learning UDFs from Multi-view Images. Different from SDFs, UDFs [9, 63, 61] are able to represent open surfaces. Recent methods [30, 35, 13, 29] design different differentiable renderes to learn UDFs through multi-view images. NeuralUDF [30] predicts the first intersection along a ray and flips the UDFs behind this point to use the differentiable renderer of NeuS [47]. NeUDF [29] proposes an inverse proportional function mapping UDF to rendering weights. NeAT [35] learns an additional validity probability net to predict the regions with open structures, while 2S-UDF [13] proposes a bell-shaped weight function that maps UDF to density, inspired by HF-NeuS [48].
The differentiable renderers introduced by these methods mainly get formulated into handcrafted equations which are biased on ray-surface intersections, sensitive to unsigned distance outliers, and not 3D aware. We resolve this issue by introducing a learning-based differentiable renderer which learns and generalizes a volume rendering prior for robustness and scalability. The ideas of learnable neural rendering frameworks are also introduced in [6, 2, 26].
@@
3 Methods
Problem Statement. Given a set of $J$ images $\{I_{j}\}_{j=1}^{J}$, we aim to infer a UDF $f_{u}$ which predicts an unsigned distance $u$ for an arbitrary 3D query $q$. We formulate the UDF as $u=f_{u}(q)$. With the learned $f_{u}$, we can extract the zero level set of $f_{u}$ as a surface using algorithms similar to the marching cubes [17, 61].
Overview. We employ a neural network to learn $f_{u}$ by minimizing rendering errors to the ground truth. We shoot rays from each view $I_{j}$, sample queries $q$ along each ray, and get unsigned distance prediction $u$ from $f_{u}$ to calculate weights $w$ for alpha blending in volume rendering. At the same time, we train a color function $c$ which predicts the color at these queries $q$ as $c=f_{c}(q)$. The accumulation of $c$ with weights $w$ along the ray produces a color at the pixel.
Current differentiable renderers [30, 13, 29, 35] transform $u$ into $w$ using handcrafted equations. Instead, we train a neural network to approximate this function $f_{w}$ in a data-driven manner, as illustrated in Fig. 4. During training, we push $f_{w}$ to produce ideal weights for rendering depth images that are as similar to the supervision $\{D_{a}^{h}\}_{a=1}^{A}$ from the $h$-th shape as possible using the ground truth UDF, and more importantly, get used to various variations of unsigned distances along a ray. During testing, we use this volume rendering prior with fixed parameters $\theta_{w}$ of $f_{w}$. We leverage $f_{w}$ to estimate an $f_{u}$ from multi-view RGB images $\{I_{j}\}$ of an unseen scene by minimizing rendering errors of RGB color through volume rendering.
Volume Rendering for UDFs. We render a UDF function $f_{u}$ with a color function $f_{c}$ into either RGB $I^{\prime}$ or depth $D^{\prime}$ images to compare with the RGB supervision $\{I_{j}\}$ or depth supervision $\{D_{j}\}$. Note that we do not use depth supervision $\{D_{j}\}$ during the UDF inference, but we include a depth supervision here to make the UDF rendering with learned priors self-contained.
From each posed view $I_{j}$, we sample some pixels and shoot rays starting at each pixel. Taking a ray $V_{k}$ from view $I_{j}$ for example, $V_{k}$ starts from the camera origin $o$ and points to a direction $r$. We hierarchically sample $N$ points along the ray $V_{k}$, where each point is sampled at $q_{n}=o+d_{n}*r$ and $d_{n}$ corresponds to the depth value of $q_{n}$ on the ray. We can transform unsigned distances $f_{u}(q_{n})$ into weights $w_{n}$ which is used for color or depth accumulation along the ray $V_{k}$ in volume rendering,
@@ 3 Methods
Although these differentiable renderers are unbiased at intersection of ray and surface and can render UDFs into images, they usually produce render errors on the boundaries on depth images, as shown by error maps in Fig. 3 (b) and (c). These errors indicate that these handcrafted equations can not render correct depth even when using the ground truth unsigned distances as supervision.
Why do these handcrafted equations produce large rendering errors? Our analysis shows that being not 3D aware plays a key role in producing these errors. These handcrafted equations merely use $M$=2 neighboring points to perceive the 3D structure when calculating the opaque density at query $q_{n}$. Such a small window makes these equations merely have a pretty small receptive field, which makes them become sensitive to unsigned distance changes, such as the weight decrease at queries sampled on a ray that is passing by an object. Moreover, to maintain some characteristics like unbiasness and occlusion awareness, these equations are strictly handcrafted, which make them extremely hard to get extended to be more 3D aware by using more neighboring points as input. Another demerit comes from the fact that all rays need to use the same equation to model the opague, which is not generalizable enough to cover various unsigned distance combinations.
@@
Learning Volume Rendering Priors. Our data-driven strategy uses ground truth meshes $\{S_{h}\}_{h=1}^{H}$ to learn the function $f_{w}$. For each shape $S_{h}$, we calculate its ground truth UDF $f_{gt}^{h}$, render $A$=100 depth images $\{D^{h}_{a}\}_{a=1}^{A}$ from randomly sampled view angles around it, and push the neural network learning $f_{w}$ to render depth images $\{\tilde{D}^{h}_{a}\}_{a=1}^{A}$ to be as similar to $\{D^{h}_{a}\}_{a=1}^{A}$ as possible. During rendering, we leverage $f_{gt}^{h}$ to provide ground truth unsigned distances at query $q_{n}$, which leaves the function $f_{w}$ as the only learning target.
@@
Specifically, along a ray $V_{k}$, we hierarchically sample $N$=128 queries $\{q_{n}\}$ to render a depth value through volume rendering using Eq.(1). We use the same sampling strategy introduced in NeUDF [29]. For each query $q_{n}$, we calculate its ground truth unsigned distance $u_{n}=f_{gt}^{h}(q_{n})$ and the ground truth unsigned distances at its $M$=30 neighboring points $\{u_{m}^{\prime}=f_{gt}^{h}(q_{m}^{\prime})\}_{m=1}^{M}$. Besides, we also use the sampling interval $\delta_{m}$ between $q_{m}^{\prime}$ and $q_{m+1}^{\prime}$ as another clue. We do not use the coordinates as a clue to pursue better generalization ability on unseen scenes with different coordinates. Therefore, we formulate the modeling of opaque density as,
$\sigma_{n}=f_{w}(\{\delta_{m}\}_{m=1}^{M},\{f_{gt}^{h}(q_{m}^{\prime})\}_{m=1}^{M}),q_{m}^{\prime}\in NN(q_{n}).$ (2)
@@
Instead of handcrafted equations, we use a neural network with 6 layers to model the function $f_{w}$. It will learn a volume rendering prior which is a prior knowledge of being a good renderer for UDFs. Obviously, it is more adaptive to different rays than handcrafted equations, and become more 3D aware with the flexibility of using a larger neighboring size. We train the network parameterized by $\theta_{w}$ by minimizing the rendering errors on depth images,
$\min_{\theta_{w}}\sum\nolimits_{h=1}^{H}\sum\nolimits_{a=1}^{A}||D_{a}^{h}-\tilde{D}^{h}_{a}||_{2}^{2}.$ (3)
@@
We do not involve RGB images in the learning of priors for better generalization ability. The improvements brought by our prior are shown in Fig. 5. We can accurately predict opaque densities with 3D awareness at arbitrary locations and robustness to unsigned distance changes.

### v2

@@ Abstract
Unsigned distance functions (UDFs) have been a vital representation for open surfaces. With different differentiable renderers, current methods are able to train neural networks to infer a UDF by minimizing the rendering errors with the UDF to the multi-view ground truth. However, these differentiable renderers are mainly handcrafted, which makes them either biased on ray-surface intersections, or sensitive to unsigned distance outliers, or not scalable to large scenes. To resolve these issues, we present a novel differentiable renderer to infer UDFs more accurately. Instead of using handcrafted equations, our differentiable renderer is a neural network which is pre-trained in a data-driven manner. It learns how to render unsigned distances into depth images, leading to a prior knowledge, dubbed volume rendering priors. To infer a UDF for an unseen scene from multiple RGB images, we generalize the learned volume rendering priors to map inferred unsigned distances in alpha blending for RGB image rendering. To reduce the bias of sampling in UDF inference, we utilize an auxiliary point sampling prior as an indicator of ray-surface intersection, and propose novel schemes towards more accurate and uniform sampling near the zero-level sets. We also propose a new strategy that leverages our pretrained volume rendering prior to serve as a general surface refiner, which can be integrated with various Gaussian reconstruction methods to optimize the Gaussian distributions and refine geometric details. Our results show that the learned volume rendering prior is unbiased, robust, scalable, 3D aware, and more importantly, easy to learn. Further experiments show that the volume rendering prior is also a general strategy to enhance other neural implicit representations such as signed distance function and occupancy. We evaluate our method on both widely used benchmarks and real scenes, and report superior performance over the state-of-the-art methods.
I Introduction
Neural implicit representations have become a dominant representation in 3D computer vision. Using coordinate-based deep neural networks, a mapping from locations to attributes at these locations like geometry [58, 53], color [13, 51], and motion [19] can be learned as an implicit representation. Signed distance function (SDF) [53] and unsigned distance function (UDF) [12] are widely used implicit representations to represent either closed surfaces [37, 9] or open surfaces [82, 21]. We can learn SDFs or UDFs from supervisions like ground truth signed or unsigned distances [5, 53], 3D point clouds [43, 61, 8, 84, 45, 44, 83, 65, 64] or multi-view images [68, 41, 26, 79, 85]. Compared to SDFs, it is a more challenging task to estimate a UDF due to the sign ambiguity and the boundary effect, especially under a multi-view setting.
@@ I Introduction
Recent methods [41, 40, 17, 46] mainly infer UDFs from multi-view images through volume rendering. Using different differentiable renderers, they can render a UDF into RGB or depth images which can be directly supervised by the ground truth images. These differentiable renderers are mainly handcrafted equations which are either biased on ray-surface intersections, or sensitive to unsigned distance outliers, or not scalable to large scale scenes. These issues make them struggle with recovering accurate geometry. Fig. 23 detail these issues by comparing error maps on depth images rendered by different differentiable renderers. We render the ground truth UDF into depth images using different renderers from 100 different view angles, and report the average rendering error on each one of 55 shapes that are randomly sampled from each one of the 55 categories in ShapeNet [6] in Fig. 2. Using the latest differentiable renderers from NeuS-UDF [62] (using UDF as input to NeuS), NeUDF [40] and NeuralUDF [41], the rendered depth images and their error maps in Fig. 3 (a)-(c) show that these issues cause large errors even using the ground truth UDF as inputs. Therefore, how to design better differentiable renderers for UDF inference from multi-view images is still a challenge.
To resolve these issues, we introduce a novel differentiable renderer for UDF inference from multi-view images through volume rendering. Instead of handcrafted equations used by the latest methods [41, 40, 17, 46], we employ a neural network to learn to become a differentiable renderer in a data-driven manner. Using UDFs and depth images obtained from meshes as ground truth, we train the neural network to map a set of unsigned distances at consecutive locations along a ray into weights for alpha blending, so that we can render depth images, and produce the rendering errors to the ground truth as a loss. We make the neural network observe different variations of unsigned distance fields during training, and learn the knowledge of volume rendering with unsigned distances by minimizing the rendering loss. The knowledge we call volume rendering prior is highly generalizable to infer UDFs from multi-view RGB images in unobserved scenes. During testing, we use the pre-trained network as a differentiable renderer for alpha blending. It renders unsigned distances inferred by a UDF network into RGB images which can be supervised by the observed RGB images. Our results in Fig. 2 and Fig. 3 (e) show that we produce the smallest rendering errors among all differentiable renderers for UDFs, which is even more accurate than NeuS-SDF [62] (rendering with ground truth SDF) in Fig. 3 (d).
@@
We originally presented our method at ECCV2024 [77], and then extend our method in more applications with novel strategies of learning volume rendering priors and unbiased sampling. Specifically, we extend the architecture of the prior network from a single window size to multiple window sizes with multi-resolution feature fusion. This design allows the network to capture detailed features at different scales according to the UDF distributions within the neighborhood for alpha blending.
We also propose two strategies to mitigate the bias in the hierarchical sampling of UDFs. First, we apply an auxiliary pre-trained point sampling prior as a ray-surface intersection indicator. The output of the indicator is used to refine the sampling densities along the ray, preventing erroneous sampling near the surfaces. Second, we delve into the inverse transform sampling process and introduce a regularization term to promote the uniform distribution of upsampled points. These approaches significantly reduce the bias and enhance the robustness of the UDF hierarchical sampling.
We additionally integrate our pretrained volume rendering priors into Gaussian reconstruction methods to serve as a general surface refiner. Taking advantage of dense queries around the UDF zero-level set, we push the Gaussian centers toward the direction of increasing accumulated rendering weights along the rays. This process facilitates the inference of accurate surfaces from sparse Gaussian distributions and significantly improves the reconstruction quality of the baseline methods.
Extensive experiments in our evaluations show that the learned volume rendering priors are unbiased, robust, scalable, 3D aware, and more importantly, easy to learn. We conduct evaluations in both widely used benchmarks and real scenes, and report superior performance over the state-of-the-art methods. We also generalize the volume rendering prior to other neural implicit representations, including SDF and occupancy, to showcase the generality and effectiveness of our volume rendering priors. Our contributions are listed below,
@@
  • We propose two novel strategies to reduce the bias of hierarchical sampling by leveraging the volume rendering prior and optimizing the inverse sampling process.
  • We extend our approach to Gaussian Splatting and propose a general strategy that consistently improves the performance of various Gaussian based reconstructions.
@@
II Related Work
II-A Multi-view 3D reconstruction
@@
Multi-view 3D reconstruction aims to reconstruct 3D shapes from calibrated images captured from overlapping viewpoints. The key idea is to leverage the consistency of features across different views to infer the geometry. MVSNet [67] is the first to introduce the deep learning-based idea into traditional MVS methods. Following studies explore the potential of MVSNet in different aspects, such as training speed [63], memory consumption [66] and network structure [18]. These techniques produce depth maps or 3D point clouds. To obtain meshes as final 3D representations, additional procedures such as TSDF-fusion [14] or classic surface reconstruction [29] methods are used, which is complex and not intuitive.
II-B Learning SDFs from Multi-view Images
@@
Instead of 3D point clouds estimated by MVS methods, recent methods [62, 68] directly estimate SDFs through volume rendering from multi-view images for continuous surface representations. The widely used strategy is to render the estimated SDF into RGB images [15, 28] or depth images [72, 60, 3] which can be supervised by the ground truth images. The key to make the whole procedure differentiable is various differentiable renderers [62, 3, 49] which transform signed distances into weights for alpha blending during rendering. Some methods modify the rendering equations to use more 2D supervisions like normal maps [59, 75], detected planes [88, 22], and segmentation maps [32, 80] to pursue higher reconstruction efficiency. Similarly, some methods [91, 90] jointly estimate camera poses and geometry in the context of SLAM. Recent efforts have attempted to recover continuous SDF fields from discrete 3D Gaussians [30]. These methods typically combine 3DGS with neural volume rendering [11, 70, 36] or differentiable pulling operations [76, 50, 39], accompanied by a series of regularization constraints [20, 25, 10, 78]. However, the SDFs that these methods aim to learn are only for closed surfaces, which is limited to represent open surfaces.
II-C Learning UDFs from Multi-view Images
@@
Different from SDFs, UDFs [12, 86, 82, 38, 74, 39] are able to represent open surfaces. Recent methods [41, 46, 17, 40] design different differentiable renderers to learn UDFs through multi-view images. NeuralUDF [41] predicts the first intersection along a ray and flips the UDFs behind this point to use the differentiable renderer of NeuS [62]. NeUDF [40] proposes an inverse proportional function mapping UDF to rendering weights. NeAT [46] learns an additional validity probability net to predict the regions with open structures, while 2S-UDF [17] proposes a bell-shaped weight function that maps UDF to density in a two-stage manner. The differentiable renderers introduced by these methods mainly get formulated into handcrafted equations which are biased on ray-surface intersections, sensitive to unsigned distance outliers, and not 3D aware. We resolve this issue by introducing a learning-based differentiable renderer which learns and generalizes a volume rendering prior for robustness and scalability. The ideas of learnable neural rendering frameworks are also introduced in [7, 2, 33].
Hierarchical Volume Sampling. NeRF [48] proposes a two-stage hierarchical volume sampling (HVS) strategy to increase the rendering efficiency. Subsequent works optimize the positions of sampling rays and sampling points by employing more comprehensive theoretical models [35], introducing auxiliary networks [2, 33], incorporating rendering results [79, 57], and improving ray representations [4]. However, these methods are mostly tailored for density-based volume rendering [48, 4] and are not applicable to neural implicit function-based volume rendering processes [62, 41]. NeuS [62] proposes an unbiased sampling strategy for SDF-based volume rendering, but this design becomes more complex when applied to UDFs, as UDFs lack an explicit distinction of zero-level set. Building upon the sampling approach of NeUDF [40], we introduce an auxiliary point sampling prior and optimize the inverse transform sampling to significantly mitigate the bias in the hierarchical sampling process.
@@
III Method
III-A Overview
Given a set of $J$ images $\{I_{j}\}_{j=1}^{J}$, we aim to infer a UDF $f_{u}$ which predicts an unsigned distance $u$ for an arbitrary 3D query $q$. We formulate the UDF as $u=f_{u}(q)$. With the learned $f_{u}$, we can extract the zero level set of $f_{u}$ as a surface using algorithms similar to the marching cubes [21, 82].
We employ a neural network to learn $f_{u}$ by minimizing rendering errors to the ground truth. We shoot rays from each view $I_{j}$, sample queries $q$ along each ray, and get unsigned distance prediction $u$ from $f_{u}$ to calculate weights $w$ for alpha blending in volume rendering. At the same time, we train a color function $c$ which predicts the color at these queries $q$ as $c=f_{c}(q)$. The accumulation of $c$ with weights $w$ along the ray produces a color at the pixel.
Current differentiable renderers [41, 17, 40, 46] transform $u$ into $w$ using handcrafted equations. Instead, we train a neural network to approximate this function $f_{w}$ in a data-driven manner, as illustrated in Fig. 4. During training, we push $f_{w}$ to produce ideal weights for rendering depth images that are as similar to the supervision $\{D_{a}^{h}\}_{a=1}^{A}$ from the $h$-th shape as possible using the ground truth UDF, and more importantly, get used to various variations of unsigned distances along a ray. During testing, we use this volume rendering prior with fixed parameters $\theta_{w}$ of $f_{w}$. We leverage $f_{w}$ to estimate an $f_{u}$ from multi-view RGB images $\{I_{j}\}_{j=1}^{J}$ of an unseen scene by minimizing rendering errors of RGB color through volume rendering.
Figure 5: Comparisons of different differentiable renderer structures. Existing methods use handcrafted equations to convert UDFs to opaque density. We extend the single-resolution MLP of VRPrior to multi-resolution MLPs, which further enhance the robustness and 3D awareness of volume rendering priors in neighborhood.
III-B Volume Rendering for UDFs
We render a UDF function $f_{u}$ with a color function $f_{c}$ into either RGB $I^{\prime}$ or depth $D^{\prime}$ images to compare with the RGB supervision $\{I_{j}\}_{j=1}^{J}$ or depth supervision $\{D_{j}\}$. From each posed view $I_{j}$, we sample some pixels and shoot rays starting at each pixel. Taking a ray $V_{k}$ from view $I_{j}$ for example, $V_{k}$ starts from the camera origin $o$ and points to a direction $r$. We hierarchically sample $N$ points along the ray $V_{k}$, where each point is sampled at $q_{n}=o+d_{n}*r$ and $d_{n}$ corresponds to the depth value of $q_{n}$ on the ray. We can transform unsigned distances $f_{u}(q_{n})$ into weights $w_{n}$ which is used for color or depth accumulation along the ray $V_{k}$ in volume rendering,
@@ III-B Volume Rendering for UDFs
Although these differentiable renderers can render UDFs into images, they are usually biased at intersection of ray and surface and produce rendering errors on the boundaries on depth images, as shown by error maps in Fig. 3 (b) and (c). These errors indicate that these handcrafted equations can not render correct depth even when using the ground truth unsigned distances as supervision.
Why do these handcrafted equations produce large rendering errors? Our analysis shows that being not 3D aware plays a key role in producing these errors. These handcrafted equations merely use MM=2 neighboring points to perceive the 3D structure when calculating the opaque density at query $q_{n}$, as shown in Fig. 5 (a). Such a small window makes these equations merely have a pretty small receptive field, which makes them become sensitive to unsigned distance changes, such as the weight decrease at queries sampled on a ray that is passing by an object. Moreover, to approximate some characteristics like unbiasedness and occlusion awareness, these equations are strictly handcrafted, which make them extremely hard to get extended to be more 3D aware by using more neighboring points as input. Another demerit comes from the fact that all rays need to use the same equation to model the opaque, which is not generalizable enough to cover various unsigned distance combinations.
@@ III-C Learning Volume Rendering Priors
Our data-driven strategy uses ground truth meshes $\{S_{h}\}_{h=1}^{H}$ to learn the function $f_{w}$. For each shape $S_{h}$, we calculate its ground truth UDF $f_{u_{gt}}^{h}$, render AA=100 ground truth depth images $\{D^{h}_{a}\}_{a=1}^{A}$ from randomly sampled view angles around it, and push the neural network learning $f_{w}$ to render depth images $\{\tilde{D}^{h}_{a}\}_{a=1}^{A}$ to be as similar to $\{D^{h}_{a}\}_{a=1}^{A}$ as possible. During rendering, we leverage $f_{u_{gt}}^{h}$ to provide ground truth unsigned distances at query $q_{n}$, which leaves the function $f_{w}$ as the only learning target.
@@
Specifically, along a ray $V_{k}$, we hierarchically sample NN=128 queries $\{q_{n}\}$ to render a depth value through volume rendering using Eq. (1). We employ a sampling strategy detailed in Sec. III-E. For each query $q_{n}$, we calculate its ground truth unsigned distance $u_{n}\!=\!f_{u_{gt}}^{h}(q_{n})$ and the ground truth unsigned distances at its neighboring points $\{u_{m}^{\prime}\!=\!f_{u_{gt}}^{h}(q_{m}^{\prime})\}_{m=1}^{M}$. Besides, we also use the sampling interval $\delta_{m}$ between $q_{m}^{\prime}$ and $q_{m+1}^{\prime}$ as another clue. Therefore, we formulate the modeling of opaque density as,
$\sigma_{n}=f_{w}(\{\delta_{m}\}_{m=1}^{M},\{f_{u_{gt}}^{h}(q_{m}^{\prime})\}_{m=1}^{M}),q_{m}^{\prime}\in NN(q_{n}).$ (2)
@@
Instead of handcrafted equations, we use a neural network to model the function $f_{w}$. It will learn a volume rendering prior which is a prior knowledge of being a good renderer for UDFs. Obviously, it is more adaptive to different rays than handcrafted equations, and become more 3D aware with the flexibility of using a larger neighboring size.
Figure 7: Comparison between VRPrior and our method. Our multi-resolution windows enable significantly more accurate transformations in handling different UDF variants, such as large near-surface regions or sharp surface corners.
To achieve this, we first divide the neighboring points into three groups based on their distances to $q_{n}$, as illustrated in Fig. 5 (c), where each group includes $M_{1}$=10, $M_{2}$=20 and $M_{3}$=30 points. For each group, we concatenate all of the unsigned distances $\{u_{m}^{\prime}\}$ and sampling intervals $\{\delta_{m}\}$, and feed them into three separate MLPs $f_{w_{1}}$ with non-shared parameters to extract three features. The features are then added together and fed into another MLP $f_{w_{2}}$ with skip connection to predict the opaque density $\sigma_{n}$ at $q_{n}$. Unlike our previous work [77] that uses a fixed single window size as shown in Fig. 5 (b), we extract features from neighborhoods with different window sizes and fuse them to produce the final result. We find that the proposed multi-resolution MLPs further enhance the generalization ability of our volume rendering priors when dealing with UDF distributions with different patterns. As highlighted in Fig. 7, when encountering scenarios such as a ray slowly passing through a surface (left in the figure) or traversing a sharp object corner (right in the figure), VRPrior [77] shows degraded performance. In contrast, our method adaptively infers the opaque densities by aggregating features from multi-scale window sizes, ensuring greater flexibility and robustness. We train all involved networks parameterized by $\theta_{w}$ by minimizing the rendering errors on depth images,
[…]
@@
Note that we do not employ RGB images in the learning of priors, because our goal at this stage is to learn a $f_{w}$ to map UDFs to opaque densities for alpha blending. The additional color network is unnecessary and would reduce the generalization ability of the learned prior to unseen cases. The improvements brought by our prior are shown in Fig. 6. We can accurately predict opaque densities with 3D awareness at arbitrary locations and robustness to unsigned distance changes.
Figure 8: Visualization of opaque density and rendering weight distributions of stage 1 and stage 2. Stage 1 perceives a broad range of potential surface locations, while stage 2 concentrates the rendering weights on the first true surface.
III-D Generalizing Volume Rendering Priors